Calculate the mass in g of potassium chloride and volume in dm3 of oxygen gas produced at S.T.P when 100g potassium chlorate is decomposed
1
Expert's answer
2021-08-18T13:28:01-0400

2KClO3….>2KCl+3O2


No of Moles of KClO3=mass/RMM=4.72/122.5=0.0385 Moles


No of Moles of oxygen=3×0.0385/2=0.058


Mass of oxygen=no.of Moles×RAM=0.058×32=1.85g


100g =70 litres


1.85g=?


70×1.85/1000=0.195 litres